OpenStudy (anonymous):

Can someone explain to me how to solve this? Find the length of an arc of 40° in a circle with an 8 inch radius.

OpenStudy (jh3power):

\[2\pi \times 8 \times \frac{40}{360}\]

OpenStudy (anonymous):

\[\frac{ 16 }{ 9 }\]

OpenStudy (anonymous):

i forgot the pi symbol but 16/9 (pi)

OpenStudy (jh3power):

absolutely

OpenStudy (anonymous):

thanks for the help